# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v2

@@ Abstract
This article proves the following theorem, first enunciated by Roger Penrose about 70 years ago: In $\mathbb{R}P^{2}$, if regular conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists a unique eighth conic such that the completed cube satisfies (i) and (ii). The proof is based on the following analogous theorem, which is also proved: In $\mathbb{R}P^{3}$, if regular quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists a unique eighth quadric such that the completed cube satisfies (i) and (ii).
@@ 1 Introduction
This article shows a “final” generalization of this progression of projective geometric theorems, first stated by Roger Penrose, the last aurthor of this article, in around 1950 during his college years. We refer to it Penrose’s eight-conic theorem (Figure 4):
If conics are assigned to seven of the vertices of a combinatorial cube such that conics connected by an edge are in double contact, then (upon a mild consistency condition on the cube face) there exists a unique eighth conic which completes the cube.
Namely, if three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ each double contact with a fourth $\mathbf{{S}}_{0}$, and another three conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ are respectively in double contact with $\mathbf{{T}}^{2},\mathbf{{T}}^{3}$, with $\mathbf{{T}}^{3},\mathbf{{T}}^{1}$, and with $\mathbf{{T}}^{3}\mathbf{{T}}^{1}$ (chosen in a consistent branch of double contact family), then there exists a unique conic $\mathbf{{T}}^{0}$ that has double contact with $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$.
@@
If quadrics in 3D are assigned to seven of the vertices of a combinatorial cube such that quadrics connected by an edge are in ring contact, then (upon a mild consistency condition on the cube face) there exists a unique eighth quadric which completes the cube.
@@
The article is structured as follows. Section 2 provides the necessary background on conics and quadrics. In Section 3, we present the main theorems: the eight-conic theorem and its 3D counterpart, the eight-quadric theorem. The proof of the eight-quadric theorem is given in Section 3.3, with the eight-conic theorem derived as a corollary through an extrusion process explained in Section 4. Section 5 illustrates examples of the eight-quadric configuration, showing how its 2D slices or projections yield classical geometric theorems in the plane. The final sections of the article provide references to related works and future outlooks.
@@ 2.2.2 Ring contact
refer to its polar point as the ring point. By the same reasoning as in the conic case, we have
@@
The polarities on two quadrics in ring contact act the same when restricted to a mapping between the points of the ring plane and the planes of the ring point. Conversely, if two polarities coincide on a plane, then the corresponding quadrics are in ring contact with respect to this plane.
@@ 3 The main theorems
In $\mathbb{R}P^{2}$, if regular conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists a unique eighth conic such that the completed cube satisfies (i) and (ii).
@@
In $\mathbb{R}P^{3}$, if regular quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists a unique eighth quadric such that the completed cube satisfies (i) and (ii).
@@ 3.1 Notation and terminology
Both the 8-conic and 8-quadric theorems assume that a regular Penrose 7-configuration is given and assert a unique completion to an 8-configuration. Note that this eighth conic (or quadric) is not necessarily regular (see Thm. 3.3).
@@
Next, introduce notation for the vertices and faces of the cube. Consult Fig. 5, left. Decompose the cube vertices into two interlocked tetrahedra. Label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$, the second is $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $F^{ij}_{km}$ where $ijkm$ is a permutation of $0123$.
@@
The face figure. Consider two conics corresponding to opposite nodes in a face of the cube, e.g. $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+pp^{T}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+qq^{T}$ with $p,q$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Hence $\mathbf{{T}}^{1}-\mathbf{{T}}^{3}=pp^{T}-qq^{T}=\frac{1}{2}(de^{T}+ed^{T})$ with $d=p+q$ and $e=p-q$ being a line pair of common chords in the pencil spanned by $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$. Thus the chords of contact meet at the point of intersection of a pair of common chords $e,d$. Since there are at most 3 such pairs of common chords the set of conics being in double contact with $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$ splits into at most 3 families, such that condition (ii) holds automatically for all the conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. All these conics can explicitly be written in terms of $d,e$ and $\mathbf{{S}}_{0}$ (see Figure 5 and [Sal17, p.262]). We call this common point the face point of the face.
@@ 3.2 Penrose’s 8-conic theorem
Apply Lemma 3.1 to the given regular Penrose 7-conic configuration to obtain an extrusion to a regular Penrose 7-quadric configuration. Then apply Thm. 3.2 to complete this to a Penrose 8-quadric configuration. The intersection of the plane of the original 7-conic configuration with the complete 8-quadric configuration is the desired regular Penrose 8-conic configuration. ∎
Remark. Due to circumstances explained in the next section, the proof presented here applies when the face axes of the Penrose 7-configurations are distinct. The only other possibility, that they are identical, will be handled using other techniques in [ACG+].
@@ 3.3 Penrose’s 8-quadric theorem
These will both be nonzero since otherwise, if we had $P^{n}_{i}=P^{n}_{j}$ or $p^{n}_{i}=p^{n}_{j}$, $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ would have to be in ring contact along the common conic of intersection with $\mathbf{{T}}^{n}$, contradicting Lemma
@@
In a regular Penrose 7-quadric configuration all ring planes and face axes are concurrent in a point $O$ or in an axis $\Omega$. Dually, all ring points and face spears are concurrent in a plane $o$ or in a spear $\omega$. The face axes are concurrent in a point exactly when the face spears are concurrent in a plane.
@@
We prove first the existence of the point $O$ in the case that the three face axes $\Lambda^{ij}_{0k}$ of the faces adjacent to $\mathbf{{S}}_{0}$ are not all the same. In this case, they are in fact all distinct because if $\Lambda^{23}_{01}=\Lambda^{13}_{02}$ then it is the shared axis of $p_{0}^{2}$ and $p_{0}^{1}$ and thus also identical with $\Lambda^{12}_{03}$.
If these 3 axes lie in a plane, then that plane must be identical to the three ring planes $p^{j}_{0}$ but by the regularity of $\mathbf{{S}}_{0}$ this would imply that $\mathbf{{T}}^{i}$ is in ring contact to $\mathbf{{T}}^{j}$ for $i,j=1,2,3$ contradicting Lemma 3.2.
Hence the three axes do not lie in a plane, and the ring planes are therefore distinct, and have a common point $O$, which is therefore also a common point of the axes.
Apply the same argument to the other complete vertices $\mathbf{{T}}^{i}$ to conclude that the axes of the faces meeting at $\mathbf{{T}}^{i}$ have a common point $O^{\prime}$. We show that $O^{\prime}=O$. Indeed, two of these axes belong to the group of 3 axes that were shown above to meet in $O$. So $O^{\prime}=O$. Since this applies to all 3 incomplete faces, all 6 axes are concurrent in $O$.
We now handle the case that all three axes $\Lambda^{ij}_{0k}$ are the same, call it $\Omega$. By reasoning similar to the above that showed that the remaining three axes are distinct, we can show that that remaining three axes also have to be equal to $\Omega$, and consequently that all ring planes lie in the plane pencil through $\Omega$.
By regularity and duality the corresponding result can be obtained for the ring points and the face spears, as stated in the theorem. By Lemma 2.2, the face axes are all distinct if and only if the face spears are all distinct.
@@
First, we assume that two of the face axes of the Penrose 7-quadric configuration are distinct. The alternative, that they are all the same, will be handled in [ACG+].
@@
To prove the theorem it suffices to construct a unique quadric $\mathbf{{T}}^{0}$ such that: for each $j\in\{1,2,3\}$, (i) $\mathbf{{T}}^{0}$ is in ring contact to $\mathbf{{S}}_{j}$; and (ii) all 4 ring planes for the face $F^{0j}_{km}$ have a common axis.
We proceed now to construct such a quadric under the assumption that the 3 axes of the complete faces of the 7-configuration are not all identical.
@@
We now show condition (i) is also fulfilled. The polarity of $\mathbf{{R}}^{0}_{i}$ in the plane $p^{0}_{i}$ maps a point $P$ to the line of intersection of $p^{0}_{i}$ and the polar plane of $P$ with respect to any quadric containing $\mathbf{{R}}^{0}_{i}$: $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ in particular. By equations (10) and (11), and Lemma 2.2, the polarities on $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$ and $\mathbf{{T}}^{0}$ all behave the same way when restricted to points on $\Lambda^{0n}_{jk}$ and planes through $\lambda^{0n}_{jk}$. Combining these two facts and applying them to the cases $i=j$ and $i=k$, any point on $\Lambda^{0n}_{jk}$ is mapped under the respective polarities of $\mathbf{{R}}^{0}_{j}$ and $\mathbf{{R}}^{0}_{k}$ to two lines which together span the polar plane of $P$ with respect to any of the quadrics $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$, $\mathbf{{T}}^{0}$. Finally, by considering the same scenario with $\Lambda^{0k}_{jn}$ in place of $\Lambda^{0n}_{jk}$, and appealling to the definitions of $P^{0}_{j}$ and $p^{0}_{j}$ in equations (12) and (13), we conclude that the polarities on $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{0}$ agree when restricted to planes through $P^{0}_{j}$ and points on $p^{0}_{j}$ so $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{j}$ are in ring contact by Lemma 2.2. ∎
@@
3.3.1 Discussion of degenerate completions
We remark that by equation (29) the $\lambda^{0n}_{ij}$ meet in a point exactly when the $x_{n}$ share a common line, i.e., when the resulting matrix $D$ is degenerate. It must, furthermore, have rank $3$ because the $\lambda^{0n}_{ij}$ do not coincide and $o$ is not incident to $O$ and thus is independent of the $x_{n}$. Therefore in case (+)(+-)( + - ), if the $\lambda^{0n}_{ij}$ are dependent, $\mathbf{{T}}^{0}$ is a cone. In case (+)(-+)( - + ), we dualize and obtain a conic. The final point-wise quadric in ring contact is the plane of this conic, doubled. In case ()(--)( - - ), the final point-wise quadric is again a double plane but now it does not arise as the plane of a conic but rather as a contact element together with a pair of lines: the lines of $X$-contact (note however that the $\mathbf{{S}}_{j}$ are not in $X$-contact with each other: as discussed in, e.g. [BO20], there are several different contact relations for which two quadrics may meet in a pair of lines, $X$-contact being the special case where this is the entirety of the intersection; in the present case, this will hold for each $\mathbf{{S}}_{j}$ with the double plane $\mathbf{{T}}^{0}$ but not for the $\mathbf{{S}}_{j}$ with each other). From this it is clear that the full information of the quadrics is not contained in the point-wise or plane-wise description alone but requires knowledge of both (as well as possibly lines). However, in the present paper we leave the precise formulation somewhat one-sided, leaving a more principled exposition to future work.

### v3

@@ Abstract
This article proves the following theorem, first enunciated by Roger Penrose about 70 years ago: In $\mathbb{R}P^{2}$, if regular conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists an eighth conic such that the completed cube satisfies (i) and (ii). This conic is unique if not all the common points of condition (ii) are the same. The proof is based on the following analogous theorem, which is also proved: In $\mathbb{R}P^{3}$, if regular quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists an eighth quadric such that the completed cube satisfies (i) and (ii). This quadric is unique if not all the common axes of condition (ii) are the same.
@@ 1 Introduction
This article shows a “final” generalization of this progression of projective geometric theorems, first stated by Roger Penrose, the last author of this article, in around 1950 during his college years. We refer to it Penrose’s eight-conic theorem (Figure 4):
If conics are assigned to seven of the vertices of a combinatorial cube such that conics connected by an edge are in double contact, then (upon a mild consistency condition on the cube face) there exists an eighth conic which completes the cube.
Namely, if three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ each in double contact with a fourth $\mathbf{{S}}_{0}$, and another three conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ are respectively in double contact with $\mathbf{{T}}^{2},\mathbf{{T}}^{3}$, with $\mathbf{{T}}^{3},\mathbf{{T}}^{1}$, and with $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$ (chosen in a consistent branch of a double contact family), then there exists a conic $\mathbf{{T}}^{0}$ that has double contact with $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$.
@@
If quadrics in 3D are assigned to seven of the vertices of a combinatorial cube such that quadrics connected by an edge are in ring contact, then (upon a mild consistency condition on the cube face) there exists an eighth quadric which completes the cube.
@@
The article is structured as follows. Section 2 provides the necessary background on conics and quadrics. In Section 3, we present the main theorems: the eight-conic theorem and its 3D counterpart, the eight-quadric theorem. The proof of the eight-quadric theorem is given in Section 3.3, with the eight-conic theorem derived as a corollary through an extrusion process explained in Section 4. Section 5 illustrates examples of the eight-quadric configuration, showing how its 2D slices or projections yield classical theorems in the projective plane. The final sections of the article provide references to related works and future outlooks.
@@ 2.2.2 Ring contact
refer to its polar point as the ring points. By the same reasoning as in the conic case, we have
@@
The polarities on two quadrics in ring contact act the same when restricted to a mapping between the points of the ring plane and the planes of the ring points. Conversely, if two polarities coincide on a plane, then the corresponding quadrics are in ring contact with respect to this plane.
@@ 3 The main theorems
In $\mathbb{R}P^{2}$, if regular conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists an eighth conic such that the completed cube satisfies (i) and (ii). This conic is unique if not all the common points of condition (ii) are the same.
@@
In $\mathbb{R}P^{3}$, if regular quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists an eighth quadric such that the completed cube satisfies (i) and (ii). This quadric is unique if not all the common axes of condition (ii) are the same.
@@ 3.1 Notation and terminology
Both the 8-conic and 8-quadric theorems assume that a regular Penrose 7-configuration is given and assert a unique completion to an 8-configuration. Note that this eighth conic (or quadric) is not necessarily unique or regular (see Thm. 3.3).
@@
Next, we introduce notation for the vertices and faces of the cube. Consult Fig. 5, left. We decompose the cube vertices into two interlocked tetrahedra, and label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$ and the second $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $F^{ij}_{km}$ where $ijkm$ is a permutation of $0123$.
In condition (ii) of the 8-conic theorem, we call the common point of the chords of contact on a face the face point of the face; for condition (ii) of the 8-quadric theorem we call the common axis of the ring planes the face axis of the face.
@@
The face figure of a Penrose conic configuration. Consider two conics corresponding to opposite nodes in a face of the cube, e.g. $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+pp^{T}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+qq^{T}$ with $p,q$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Hence $\mathbf{{T}}^{1}-\mathbf{{T}}^{3}=pp^{T}-qq^{T}=\frac{1}{2}(de^{T}+ed^{T})$ with $d=p+q$ and $e=p-q$ being a line pair of common chords in the pencil spanned by $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$. Thus the chords of contact meet at the point of intersection of a pair of common chords $e,d$. Since there are at most 3 such pairs of common chords the set of conics being in double contact with $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$ splits into at most 3 families, such that condition (ii) holds automatically for all the conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. All these conics can explicitly be written in terms of $d,e$ and $\mathbf{{S}}_{0}$ (see Figure 5 and [Sal17, p.262]).
@@ 3.2 Penrose’s 8-conic theorem
Apply Lemma 3.1 to the given regular Penrose 7-conic configuration to obtain an extrusion to a regular Penrose 7-quadric configuration. Then apply Thm. 3.2 to complete this to a Penrose 8-quadric configuration. The intersection of the plane of the original 7-conic configuration with the complete 8-quadric configuration is the desired regular Penrose 8-conic configuration. Face points of the conic configuration are distinct exactly when the corresponding face axes of the quadric configuration are distinct. Hence the eighth conic is unique exactly when the face points are not all the same. ∎
Remark. Due to factors explained in the next section, the proof presented below for the 8-quadric theorem does not cover the case that all the face axes, resp., points, are identical, hence the 8-conic theorem proof above does not cover the case that all chords of contact of the configuration pass through a single point.
@@ 3.3 Penrose’s 8-quadric theorem
These will both be nonzero since otherwise, if we had $P^{n}_{i}=P^{n}_{j}$ or $p^{n}_{i}=p^{n}_{j}$, $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ would have to be in ring contact along the common conic of intersection with $\mathbf{{T}}^{n}$, contradicting Lemma 3.2.
@@
For a regular Penrose 7-quadric configuration: If two of the face axes at $\mathbf{{S}}_{0}$ are distinct, then at each complete vertex, all three face axes are distinct. In this case, all ring planes and face axes are concurrent in a point $O$. Otherwise there is a single face axis $\Omega$ common to all ring planes. Dually, all ring points and face spears are concurrent in a plane $o$ or there is a single face spear $\omega$ common to all ring points. $O$ (resp., $\Omega$) and $o$ (resp., $\omega$) are a polar pair with respect to all quadrics of the configuration.
@@
Assume $\Lambda^{23}_{01}=\Lambda^{13}_{02}$. Since each axis contains two of the ring planes at $\mathbf{{S}}_{0}$, all three of these ring planes pass through this axis so that it is also identical to $\Lambda^{12}_{03}$.
This contradicts the assumption that two of the three face axes $\Lambda^{ij}_{0k}$ are distinct. Hence all three are distinct if any two are.
We show now that the $\Lambda^{ij}_{0k}$ cannot all lie in a plane. Indeed, if they do, then that plane must be identical to the three ring planes $p^{j}_{0}$ but by the regularity of $\mathbf{{S}}_{0}$ this would imply that $\mathbf{{T}}^{i}$ is in ring contact to $\mathbf{{T}}^{j}$ for $i,j=1,2,3$ contradicting Lemma 3.2.
Hence the three axes at $\mathbf{{S}}_{0}$ are distinct and do not lie in a plane, and the ring planes are therefore distinct, and have a common point $O$, which is therefore also a common point of the axes.
Apply the same argument to the other complete vertices $\mathbf{{T}}^{i}$ to conclude that the axes of the faces meeting at $\mathbf{{T}}^{i}$ have a common point $O^{\prime}$. We show that $O^{\prime}=O$. Indeed, two of these axes belong to $\{\Lambda^{ij}_{0k}\}$, that were shown above to meet in $O$. So $O^{\prime}=O$. Since this applies to all 3 incomplete faces, all 6 axes are concurrent in $O$.
We now consider the case that two of the axes at $\mathbf{{S}}_{0}$ are the same. Then by the above, all three axes $\Lambda^{ij}_{0k}$ are the same, call it $\Omega$. By reasoning similar to the above, we can propagate this condition to the three remaining face axes, to show they also have to be equal to $\Omega$, and consequently that all ring planes pass through $\Omega$.
By regularity and duality the analogous result can be obtained for the ring points and the face spears. Furthermore, by Lemma 2.2, $O$ (resp., $\Omega$) and $o$ (resp., $\omega$) are a polar pair with respect to all quadrics of the configuration.
@@ 3.3.1 The proof
First, we assume that two of the face axes at $\mathbf{{S}}_{0}$ are distinct. By Lemma 3.3, the alternative is that all six face axes are the same. This case will be handled using other methods in [ACG+].
@@
Under the assumption that two of the faces axes at $\mathbf{{S}}_{0}$ are distinct, it suffices to construct a unique quadric $\mathbf{{T}}^{0}$ such that: for each $j\in\{1,2,3\}$, (i) $\mathbf{{T}}^{0}$ is in ring contact to $\mathbf{{S}}_{j}$; and (ii) all 4 ring planes for the face $F^{0j}_{km}$ have a common axis.
@@
We now show condition (i) is also fulfilled. The polarity of $\mathbf{{R}}^{0}_{i}$ in the plane $p^{0}_{i}$ maps a point $P$ to the line of intersection of $p^{0}_{i}$ and the polar plane of $P$ with respect to any quadric containing $\mathbf{{R}}^{0}_{i}$: $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ in particular. By equations (10) and (11), and Lemma 2.2, the polarities on $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$ and $\mathbf{{T}}^{0}$ all behave the same way when restricted to points on $\Lambda^{0n}_{jk}$ and planes through $\lambda^{0n}_{jk}$. Combining these two facts and applying them to the cases $i=j$ and $i=k$, any point on $\Lambda^{0n}_{jk}$ is mapped under the respective polarities of $\mathbf{{R}}^{0}_{j}$ and $\mathbf{{R}}^{0}_{k}$ to two lines which together span the polar plane of $P$ with respect to any of the quadrics $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$, $\mathbf{{T}}^{0}$. Finally, by considering the same scenario with $\Lambda^{0k}_{jn}$ in place of $\Lambda^{0n}_{jk}$, and appealing to the definitions of $P^{0}_{j}$ and $p^{0}_{j}$ in equations (12) and (13), we conclude that the polarities on $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{0}$ agree when restricted to planes through $P^{0}_{j}$ and points on $p^{0}_{j}$ so $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{j}$ are in ring contact by Lemma 2.2. ∎
@@
3.3.2 When the eighth quadric is not regular
The case (++)(++)( + + ), in which both triples of axes and spears are linearly independent, clearly produces regular matrices, hence the eighth quadric is regular. This section considers what kinds of degenerate quadrics arise in the other cases. We remark that by equation (29) the $\lambda^{0n}_{ij}$ meet in a point exactly when the $x_{n}$ share a common line, i.e., when the resulting matrix $D$ is degenerate. It must, furthermore, have rank $3$ because the $\lambda^{0n}_{ij}$ do not coincide and $o$ is not incident to $O$ and thus is independent of the $x_{n}$. Therefore in case (+)(+-)( + - ), if the $\lambda^{0n}_{ij}$ are dependent, $\mathbf{{T}}^{0}$ is a cone. In case (+)(-+)( - + ), we dualize and obtain a conic. The final point-wise quadric in ring contact is the plane of this conic, doubled. In case ()(--)( - - ), the final point-wise quadric is again a double plane but now it does not arise as the plane of a conic but rather as a contact element together with a pair of lines: the lines of $X$-contact (note however that the $\mathbf{{S}}_{j}$ are not in $X$-contact with each other: as discussed in, e.g. [BO20], there are several different contact relations for which two quadrics may meet in a pair of lines, $X$-contact being the special case where this is the entirety of the intersection; in the present case, this will hold for each $\mathbf{{S}}_{j}$ with the double plane $\mathbf{{T}}^{0}$ but not for the $\mathbf{{S}}_{j}$ with each other). From this it is clear that the full information of the quadrics is not contained in the point-wise or plane-wise description alone but requires knowledge of both (as well as possibly lines). However, in the present paper we leave the precise formulation somewhat one-sided, leaving a more principled exposition to future work.